the point of intersection of the tangents to the parabola y^2=4ax at the points t1 and t2 is....

Grade:11

1 Answers

Vikas TU
7 years ago
Simply write the eqn. of tangent for parabola y^2 = 4ax in parametric form with t1 and t2 respectively.
And then solve them .
that is to say that:
t1y = x + at1^2 …..................(1)
and
t2y = x + at2^2......................(2)
solving them U would get as the point of intersection as:
(at1t2 , a(t1+t2))